Consider the following future value. (Round your answers to the nearest cent.)

$1,000 at 8% compounded annually for 7 years

(a) Find the present value that will generate the given future value.
$ ...

(b) Interpret the present value.

One would have to invest $ ... now to have the future value in the given time.
Answer #1

FV = PV*(1+r)^n

1000 = PV*(1+0.08)^7

1000 = PV*1.71382427

PV= 1000/ 1.71382427

= $583.49

Hence one would have to invest $ 583.49 now to have the future value of $1000 at the given time of 7 years at 8% interest rate.